A particle executes shm along y axis has motion described by equation y=Asin (wt)+B the amplitude of shm

Grade:12

1 Answers

Shravan Chaudhari
6 years ago
See amplitude of a periodic motion is maximum value of of its displacement from mean position in general equation y=Asin(wt) we get maximum value when t=pi/2w that is y=A . Here we want maximum of y=Asin(wt)+ B  which will be when sinwt=1 so maximum value of the given equation comes out A+B which is the amplitude of the shm
Hope u got it